# Breaking the Exploration Bottleneck: Rubric-Scaffolded Reinforcement Learning for General LLM Reasoning

### v5

@@ Abstract
Recent advances in Large Language Models (LLMs) have underscored the potential of Reinforcement Learning (RL) to facilitate the emergence of reasoning capabilities. Despite the encouraging results, a fundamental dilemma persists as RL improvement relies on learning from high-quality samples, yet the exploration for such samples remains bounded by the inherent limitations of LLMs. This, in effect, creates an undesirable cycle in which what cannot be explored cannot be learned. In this work, we propose Rubric-Scaffolded Reinforcement Learning (RuscaRL), a novel instructional scaffolding framework designed to break the exploration bottleneck for general LLM reasoning. Specifically, RuscaRL introduces checklist-style rubrics as (1) explicit scaffolding for exploration during rollout generation, where different rubrics are provided as external guidance within task instructions to steer diverse high-quality responses. This guidance is gradually decayed over time, encouraging the model to internalize the underlying reasoning patterns; (2) verifiable rewards for exploitation during model training, where we can obtain robust LLM-as-a-Judge scores using rubrics as references, enabling effective RL on general reasoning tasks. Extensive experiments demonstrate the superiority of the proposed RuscaRL across various benchmarks, effectively expanding reasoning boundaries under the Best-of-N evaluation. Notably, RuscaRL significantly boosts Qwen2.5-7B-Instruct from 23.6 to 50.3 on HealthBench-500, surpassing GPT-4.1. Furthermore, our fine-tuned variant on Qwen3-30B-A3B-Instruct achieves 61.1 on HealthBench-500, outperforming leading LLMs including OpenAI-o3. Our code is available at https://github.com/IANNXANG/RuscaRL.
##footnotetext: Corresponding author.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated tremendous potential over a wide spectrum of complex reasoning tasks, including legal analysis (Choi et al., 2021; Lai et al., 2024; Fei et al., 2023; Trautmann et al., 2022), robotic manipulation (Driess et al., 2023; Zitkovich et al., 2023; Firoozi et al., 2025; Zhou et al., 2023a), and software development (Anysphere, 2023; Fan et al., 2023; Hou et al., 2024). However, advancing the general reasoning capabilities of LLMs remains a significant challenge (Zhao et al., 2023; Huang & Chang, 2022). To address this, recent breakthroughs in Reinforcement Learning with Verifiable Rewards (RLVR), as exemplified by DeepSeek-R1 (Guo et al., 2025) and OpenAI-o3 (OpenAI, 2025a), have proven that leveraging verifiable rewards as feedback signals can successfully facilitate the emergence of sophisticated reasoning capabilities in LLMs (Lambert et al., 2024; Yang et al., 2025; Kimi et al., 2025).
Despite the encouraging results, conventional RLVR tends to be more applicable to domains with objectively verifiable answers. For instance, in areas such as mathematical proof (Ren et al., 2025; Chen et al., 2025) and code generation (Qwen, 2025; Le et al., 2022), correctness can be explicitly determined through formal proof verification or automated unit tests. In these contexts, the reward signal is clear and well-aligned with the task objective, allowing RLVR to effectively guide models toward correct solutions. Unfortunately, many real-world tasks like medical consultation (Lin et al., 2025; Singhal et al., 2023; Zhang et al., 2023) and creative writing (Wu et al., 2025c; Franceschelli & Musolesi, 2024) are inherently open-ended. These tasks often require multidimensional evaluation and lack a single, verifiable ground-truth answer. To tackle this problem, several very recent concurrent works (Arora et al., 2025; Kimi et al., 2025; Gunjal et al., 2025; Viswanathan et al., 2025; Huang et al., 2025; Dou et al., 2025) have explored rubric-based evaluation that decomposes desirable responses into checklist-style criteria (e.g., factuality, coherence, completeness). By leveraging LLM-as-a-Judge to score each criterion and aggregating results into scalar rewards, rubrics provide more stable and reliable feedback signals suitable for RLVR in open-ended domains.
Nevertheless, as shown in the left side of Figure 1, a fundamental exploration bottleneck remains as RL requires high-quality samples to improve, yet exploration for such samples remains bounded by the inherent capabilities of LLMs (Yue et al., 2025; Wu et al., 2025a; Liu et al., 2025b; Dong et al., 2025). This creates an inevitable loop where the inability to explore restricts the ability to learn. A growing line of work has attempted to enhance exploration in RLVR for LLMs (Liu et al., 2025b; a; Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025). However, these methods largely bias the policy distribution toward high-reward responses already supported by base models, rather than truly expanding its reasoning boundaries (Wu et al., 2025a). Worse still, RL itself has a natural tendency to narrow the exploration space: policy entropy gradually collapses during training, causing the model to converge toward a limited set of reasoning trajectories (Zhao et al., 2025; Yue et al., 2025; Wu et al., 2025a; Yu et al., 2025; Liu et al., 2025b). This, in turn, undermines the potential of RLVR to explore more diverse and higher-quality solutions.
@@
Figure 1: (Left) A conceptual illustration of exploration bottleneck and scaffolding. (Right) Performance comparison of different LLMs on HealthBench-500. Note that Qwen3-30B-Instruct denotes Qwen3-30B-A3B-Instruct, and Qwen3-235B-Thinking denotes Qwen3-235B-A22B-Thinking.
@@
  • Extensive experiments demonstrate that RuscaRL yields results superior to state-of-the-art counterparts. Notably, RuscaRL enables small LLMs (e.g., Qwen3-30B) to achieve performance on par with leading LLMs (e.g., OpenAI-o3) on HealthBench-500, as shown in the right side of Figure 1.
@@ 2 Related Works
Rubric-based Methods. Rubrics are structured evaluation frameworks that decompose complex assessment tasks into specific, verifiable criteria. To address general task evaluation, rubric-based evaluation approaches have emerged across medical (Arora et al., 2025), code (Pathak et al., 2025), and other domains (Fan et al., 2025; Galvan-Sosa et al., 2025; Winata et al., 2025). Building upon these frameworks, researchers apply rubrics as reward signals in RL (Kimi et al., 2025; Gunjal et al., 2025; Viswanathan et al., 2025), using LLMs as graders to provide fine-grained feedback for tasks lacking ground truth. This approach has shown promising results across LLM alignment (Dineen et al., 2025), instruction following (Viswanathan et al., 2025), and open-ended QA (Gunjal et al., 2025; Huang et al., 2025; Dou et al., 2025; Kimi et al., 2025).
Exploration in RL for LLMs. Existing RL methods face insufficient exploration in complex reasoning tasks, with policies trapped in local optima and reasoning boundaries collapsing (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b). Current solutions include prolonged training (Liu et al., 2025b; a), entropy-based exploration (Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025), and external guidance (Zhang et al., 2025a; RRV et al., 2025), but these approaches fail to break the exploration bottleneck because they either explore within the initial policy distribution or provide only coarse directional signals without structured intermediate guidance. In contrast, RuscaRL provides explicit scaffolding via checklist-style rubrics, guiding trajectories with verifiable criteria while enabling scaffolding decay for pattern internalization.
@@ 3 Preliminary
RL Algorithms for LLMs. We adopt Group Relative Policy Optimization  (GRPO) (Shao et al., 2024) as our core RL algorithm for training language models with rubric-based rewards. Unlike Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO eliminates the need for a value model by using group-based advantage estimation. For each instruction $q\sim\mathcal{D}$, where $\mathcal{D}$ denotes the distribution over the training dataset $\mathcal{D}$, GRPO samples a group of $G$ responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the policy $\pi_{\theta}$ by maximizing the following objective:
$\begin{split}\mathcal{J}_{\text{GRPO}}\left(\theta\right)&=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\\ &\quad\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\left(\rho_{i,t}(\theta)\hat{A}_{i},\text{clip}\left(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)\right],\end{split}$ (1)
@@
where $r_{i}$ is the reward for response $o_{i}$, and the advantage is normalized using the mean and standard deviation of the $G$ sampled rewards.
@@ 4 Methodology
Figure 2: Overview of the RuscaRL framework: (a) Rubric data sample with criteria and points, (b) Rubric-scaffolded reinforcement learning with reward computation and policy updates, (c) Intra-group scaffolding differentiation and inter-step scaffolding decay mechanisms.
To address the exploration bottleneck problem, we propose the RuscaRL framework, as illustrated in Figure 2. RuscaRL leverages rubrics in two complementary ways: (1) Explicit scaffolding during rollout generation, where the model generates candidate responses using rubrics as external guidance with intra-group differentiation and inter-step decay (Section 4.2); (2) Verifiable rewards during model training, where responses are assessed based on multiple criteria derived from rubrics through binary evaluation and aggregation (Section 4.3). In the following, we first introduce the basic concept of rubrics, then detail these two core components.
@@ 4.1 Rubric-based Evaluation System
4.2 Rubric-based Scaffolding Mechanism for RL Exploration
During RL training on complex reasoning tasks, models often fail to sustain effective exploration (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025b): initial randomness quickly diminishes, policy entropy collapses, and the model prematurely converges to suboptimal reasoning patterns. This collapse severely limits the discovery of diverse and high-quality solution trajectories.
To mitigate this issue, we draw inspiration from instructional scaffolding theory in educational psychology (Vygotsky & Cole, 1978). According to Vygotsky’s Zone of Proximal Development, when learners’ capabilities are insufficient, they benefit from structured guidance to bridge the gap between current ability and target performance. As competence grows, this scaffolding should be gradually withdrawn to foster independent problem-solving (Wood et al., 1976).
@@ 4.2 Rubric-based Scaffolding Mechanism for RL Exploration
Intra-Group Scaffolding Differentiation. In RL algorithms with multi-sampling, such as GRPO, computing group-relative advantages (Eq. 2) requires response diversity to avoid collapse into homogeneous samples. To this end, we assign different levels of rubric scaffolding within each group, encouraging both guided and independent exploration. Concretely, we define a group-level ratio vector $\bm{\lambda}_{group}=[\lambda_{1},\lambda_{2},\ldots,\lambda_{G}]$ where $\lambda_{i}=\frac{G-i}{G-1}$ for the $i$-th sample in the group of size $G$. This linear differentiation ensures that some samples benefit from stronger scaffolding while others are deliberately exposed to weaker guidance, thereby enhancing intra-group diversity.
@@
Integrated Scaffolding Mechanism. Finally, we combine intra-group differentiation and inter-step decay into a unified ratio vector:
$\bm{\lambda}_{S}=\lambda_{step}\left(t\right)\times\bm{\lambda}_{group}=\left[\lambda_{S,1},\lambda_{S,2},\ldots,\lambda_{S,G}\right],$ (3)
@@
4.3 Rubric-based Reward Function for RL Exploitation
To provide robust and reliable reward signals for general reasoning tasks, we design rubric-based reward functions. The multi-dimensional score vector $\mathbf{s}=[s_{1},s_{2},\ldots,s_{N}]$ obtained from the evaluation system is aggregated into a final scalar reward by directly summing all criterion scores and normalizing by the total possible score computed in Section 4.1:
@@ 5 Experiments
To demonstrate the effectiveness of the proposed RuscaRL method, we conduct experiments across multiple benchmarks spanning medical, writing, instruction following, and STEM domains. Our experiments seek to answer the following questions: (1) Does RuscaRL demonstrate consistent effectiveness across different models and tasks, and how does it compare against existing fine-tuning methods? (Section 5.2 and Appendix D.1, D.2, D.3) (2) How does RuscaRL break the exploration bottleneck in RL for LLM reasoning? (Section 5.3 and Appendix D.4) (3) What is the impact of different components in the rubric-based scaffolding mechanism? (Section 5.4 and Appendix D.5)

### v6

@@ Abstract
Recent advances in Large Language Models (LLMs) have underscored the potential of Reinforcement Learning (RL) to facilitate the emergence of reasoning capabilities. Despite the encouraging results, a fundamental dilemma persists as RL improvement relies on learning from high-quality samples, yet the exploration for such samples remains bounded by the inherent limitations of LLMs. This, in effect, creates an undesirable cycle in which what cannot be explored cannot be learned. In this work, we propose Rubric-Scaffolded Reinforcement Learning (RuscaRL), a novel instructional scaffolding framework designed to break the exploration bottleneck for general LLM reasoning. Specifically, RuscaRL introduces checklist-style rubrics as (1) explicit scaffolding for exploration during rollout generation, where different rubrics are provided as external guidance within task instructions to steer diverse high-quality responses. This guidance is gradually decayed over time, encouraging the model to internalize the underlying reasoning patterns; (2) verifiable rewards for exploitation during model training, where we can obtain robust LLM-as-a-Judge scores using rubrics as references, enabling effective RL on general reasoning tasks. Extensive experiments demonstrate the superiority of the proposed RuscaRL across various benchmarks, effectively expanding reasoning boundaries under the Best-of-N evaluation. Our code is available at https://github.com/IANNXANG/RuscaRL.
Machine Learning, ICML
1Zhejiang University, 2Li Auto Inc., 3Nanyang Technological University,
4The Chinese University of Hong Kong, 5Hangzhou City University
imzhouyang@zju.edu.cn, lisunzhu@lixiang.com, shunyu.liu@ntu.edu.sg
footnotetext: Corresponding author
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated tremendous potential over a wide spectrum of complex reasoning tasks, including legal analysis (Choi et al., 2021; Lai et al., 2024; Fei et al., 2023; Trautmann et al., 2022), robotic manipulation (Driess et al., 2023; Zitkovich et al., 2023; Firoozi et al., 2025; Zhou et al., 2023b), and software development (Chen, 2021; Fan et al., 2023; Hou et al., 2024). However, advancing the general reasoning capabilities of LLMs remains a significant challenge (Zhao et al., 2023; Huang and Chang, 2022). To address this, recent breakthroughs in Reinforcement Learning with Verifiable Rewards (RLVR), as exemplified by DeepSeek-R1 (Guo et al., 2025), have proven that leveraging verifiable rewards as feedback signals can successfully facilitate the emergence of sophisticated reasoning capabilities in LLMs (Lambert et al., 2024; Yang et al., 2025; Kimi et al., 2025).
@@
Figure 1: (Left) A conceptual illustration of exploration bottleneck and scaffolding. (Right) Performance comparison of different LLMs on HealthBench-500. Note that Qwen3-30B-Instruct denotes Qwen3-30B-A3B-Instruct-2507, and Qwen3-235B-Thinking denotes Qwen3-235B-A22B-Thinking-2507.
Despite the encouraging results, conventional RLVR tends to be more applicable to domains with objectively verifiable answers. For instance, in areas such as mathematical proof (Ren et al., 2025; Chen et al., 2025) and code generation (Qwen, 2025; Le et al., 2022), correctness can be explicitly determined through formal proof verification or automated unit tests. In these contexts, the reward signal is clear and well-aligned with the task objective, allowing RLVR to effectively guide models toward correct solutions. Unfortunately, many real-world tasks like medical consultation (Lin et al., 2025; Singhal et al., 2023; Zhang et al., 2023) and creative writing (Wu et al., 2025c; Franceschelli and Musolesi, 2024) are inherently open-ended. These tasks often require multidimensional evaluation and lack a single, verifiable ground-truth answer. To tackle this problem, several recent works (Arora et al., 2025; Kimi et al., 2025; Gunjal et al., 2025; Dou et al., 2025; Liu et al., 2025a) have explored rubric-based evaluation that decomposes desirable responses into checklist-style criteria (e.g., factuality, coherence, completeness). By leveraging LLM-as-a-Judge to score each criterion and aggregating results into scalar rewards, rubrics provide more stable and reliable feedback signals suitable for RLVR in open-ended domains.
Nevertheless, as shown in the left side of Figure 1, a fundamental exploration bottleneck remains as RL requires high-quality samples to improve, yet exploration for such samples remains bounded by the inherent capabilities of LLMs (Yue et al., 2025; Wu et al., 2025a; Liu et al., 2025c; Dong et al., 2025). This creates an inevitable loop where the inability to explore restricts the ability to learn. A growing line of work has attempted to enhance exploration in RLVR for LLMs (Liu et al., 2025c, b; Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025). However, these methods largely bias the policy distribution toward high-reward responses already supported by base models, rather than truly expanding its reasoning boundaries (Wu et al., 2025a). Worse still, RL itself has a natural tendency to narrow the exploration space: policy entropy gradually collapses during training, causing the model to converge toward a limited set of reasoning trajectories (Zhao et al., 2025; Yue et al., 2025; Wu et al., 2025a; Yu et al., 2025; Liu et al., 2025c). This, in turn, undermines the potential of RLVR to explore more diverse and higher-quality solutions.
@@
  • Extensive experiments demonstrate that RuscaRL yields highly competitive results compared to strong baseline methods. Notably, RuscaRL enables small LLMs (e.g., Qwen3-30B-A3B-Instruct-2507) to achieve performance on par with leading LLMs (e.g., OpenAI-o3) on HealthBench-500, as shown in the right side of Figure 1.
@@ 2 Related Works
Rubric-based Methods. Rubrics are structured evaluation frameworks that decompose complex assessment tasks into specific, verifiable criteria. To address general task evaluation, rubric-based evaluation approaches have emerged across medical (Arora et al., 2025), code (Pathak et al., 2025), and other domains (Fan et al., 2025; Galvan-Sosa et al., 2025; Winata et al., 2025). Building upon these frameworks, researchers apply rubrics as reward signals in RL (Kimi et al., 2025; Liu et al., 2025a; Gunjal et al., 2025), using LLMs as graders to provide fine-grained feedback for tasks lacking ground truth. This approach has shown promising results across agentic search (Liu et al., 2025a; Shao et al., 2025; Kimi et al., 2025), instruction following (Viswanathan et al., 2025; Huang et al., 2025), and medical consultation (Gunjal et al., 2025; Dou et al., 2025).
Exploration in RL for LLMs. Existing RL methods face insufficient exploration in complex reasoning tasks, with policies trapped in local optima and reasoning boundaries collapsing (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025c). Current solutions include prolonged training (Liu et al., 2025c, b), entropy-based exploration (Dong et al., 2025; Zheng et al., 2025; Lei et al., 2025; Li et al., 2025; Cheng et al., 2025), and external guidance (Zhang et al., 2025a; RRV et al., 2025), but these approaches fail to break the exploration bottleneck because they either explore within the initial policy distribution or provide only coarse directional signals without structured intermediate guidance. In contrast, RuscaRL provides rubrics as explicit scaffolding, guiding trajectories with verifiable criteria while enabling scaffolding decay for pattern internalization.
@@ 3 Preliminary
RL Algorithms for LLMs. We adopt Group Relative Policy Optimization  (GRPO) (Shao et al., 2024) as our core RL algorithm for training language models with rubric-based rewards. Unlike Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO eliminates the need for a value model by using group-based advantage estimation. For each instruction $q\sim\mathcal{D}$, where $\mathcal{D}$ denotes the distribution over the training dataset, GRPO samples a group of $G$ responses $\{o_{1},o_{2},\ldots,o_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$ and optimizes the policy $\pi_{\theta}$ by maximizing the following objective:
$\mathcal{J}_{\text{GRPO}}\left(\theta\right)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|q)}\bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}$ (1)
$\min\left(\rho_{i,t}(\theta)\hat{A}_{i},\text{clip}\left(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)\bigg],$
@@
where $r_{i}$ is the reward for response $o_{i}$, and the advantage is normalized using the mean and standard deviation of the $G$ sampled rewards within each group.
@@ 4 Methodology
Figure 2: Overview of the RuscaRL framework. (a) Rubric as Verifiable Rewards for Exploitation: checklist-style rubric criteria with associated points are used by an LLM-based grader to perform criterion-wise binary judgments and aggregate the resulting scores into a scalar reward for each response. (b) Rubric as Explicit Scaffolding for Exploration: the same rubrics are injected into instructions as external guidance during rollout generation, with intra-group scaffolding differentiation providing different levels of criteria to samples within a group and inter-step scaffolding decay gradually reducing the amount of scaffolding over training.
To address the exploration bottleneck problem, we propose the novel RuscaRL framework, as illustrated in Figure 2. RuscaRL effectively leverages rubrics in two complementary ways: (1) Explicit scaffolding during rollout generation, where the model generates candidate responses using rubrics as external guidance with intra-group differentiation and inter-step decay (Section 4.2); (2) Verifiable rewards during model training, where model responses are assessed based on multiple criteria derived from rubrics through binary evaluation and aggregation (Section 4.3). In the following, we first introduce the basic concept of rubrics, and then detail these two core components.
@@ 4.1 Rubric-based Evaluation System
4.2 Rubric-based Scaffolding for RL Exploration
During RL training on complex reasoning tasks, models often fail to sustain effective exploration (Wu et al., 2025a; Yue et al., 2025; Liu et al., 2025c): initial randomness quickly diminishes, policy entropy collapses, and the model prematurely converges to suboptimal reasoning patterns. This collapse severely limits the discovery of diverse and high-quality solution trajectories.
To mitigate this issue, we draw inspiration from instructional scaffolding theory in educational psychology (Vygotsky and Cole, 1978). According to Vygotsky’s Zone of Proximal Development, when learners’ capabilities are insufficient, they benefit from structured guidance to bridge the gap between current ability and target performance. As competence grows, this scaffolding should be gradually withdrawn to foster independent problem-solving (Wood et al., 1976).
@@ 4.2 Rubric-based Scaffolding for RL Exploration
Intra-Group Scaffolding Differentiation. In RL algorithms with multi-sampling, such as GRPO, computing group-relative advantages (Eq. 2) requires response diversity to avoid collapse into homogeneous samples. To this end, we assign different levels of rubric scaffolding within each group, encouraging both guided and independent exploration. Concretely, we define a group-level ratio vector $\boldsymbol{\lambda}_{group}=[\lambda_{1},\lambda_{2},\ldots,\lambda_{G}]$ where $\lambda_{i}=\frac{G-i}{G-1}$ for the $i$-th sample in the group of size $G$. This linear differentiation ensures that some samples benefit from stronger scaffolding while others are deliberately exposed to weaker guidance, thereby enhancing intra-group diversity.
@@
Integrated Scaffolding Mechanism. Finally, we combine the intra-group differentiation and the inter-step decay into a unified scaffolding ratio vector:
$\boldsymbol{\lambda}_{S}=\lambda_{step}\left(t\right)\times\boldsymbol{\lambda}_{group}=\left[\lambda_{S,1},\lambda_{S,2},\ldots,\lambda_{S,G}\right],$ (3)
@@
4.3 Rubric-based Reward for RL Exploitation
To provide robust and reliable reward signals for general reasoning tasks, we design the following rubric-based reward functions. The multi-dimensional score vector $\mathbf{s}=[s_{1},s_{2},\ldots,s_{N}]$ obtained from the evaluation system is then aggregated into a final scalar reward by directly summing all criterion scores and normalizing by the total possible score computed in Section 4.1:
@@ 5 Experiments
To demonstrate the effectiveness of the proposed RuscaRL method, we conduct experiments across multiple benchmarks spanning medical, writing, instruction following, and STEM domains. Our experiments seek to answer the following questions: (1) Does RuscaRL demonstrate consistent effectiveness across different models and tasks, and how does it compare against existing fine-tuning methods? (Section 5.2 and Appendix D.1D.2D.3D.4) (2) How does RuscaRL break the exploration bottleneck in RL for LLM reasoning? (Section 5.3 and Appendix D.5) (3) What is the impact of different components in the rubric-based scaffolding mechanism? (Section 5.4 and Appendix D.6)